# How digital will the future be? Analysis of prospective scenarios

### v1

@@ Appendix A Summary of studies reviewed for this paper
Ademe 2022
This study from French Agency for Ecological Transition ADEME proposes five coherent and contrasting paths leading France towards carbon neutrality by 2050.
@@
negaWatt 2021
Scenarios for France made by the négaWatt association to achieve carbon neutrality by 2050 while reducing the extraction of raw materials. This scenario is also compatible with the -55% greenhouse gas target set at European level for 2030.
@@
EU green deal 2019
This communication sets out the strategy to transform the EU into a fair and prosperous society, with a modern, resource-efficient and competitive economy where there are no net emissions of greenhouse gases in 2050 and where economic growth is decoupled from resource use.
@@
Eionet 2022
This study produces a set of imaginaries offering engaging, plausible and contrasting images of what a sustainable Europe could look like in 2050. The imaginaries were developed through a participatory process, involving European Environment Agency staff, experts from the Eionet group on foresight, and external stakeholders.
@@
Arup 2019
Study made by an international multi-disciplinary team of consultants, designers and scientists. In the study, Climate considerations come third, subordinate to economic development and societal wellbeing. Scenarios are narratives and accompanied by a story told from the perspective of a person within that world, a timeline of events, and the achievement of the 17 UN SDGs.
@@
DDC 2020
The Danish Design Center (DDC) designed the Living Futures: Scenario Kit for understanding, discussing, and shaping the future. The kit is built around a set of four alternative futures set in the year 2050 explored through narrated stories by fictive people living in them.
@@
SNBC 2020
The National Low Carbon Strategy (SNBC) describes a roadmap for France on how to steer its climate change mitigation policy. The SNBC is based on a reference scenario developed through a modelling exercise also used in the Multi-annual Energy Programming. The scenario details the public policy measures, in addition to those already in place, which will allow France to meet its short-, medium- and long-term climate and energy objectives as best it can.
@@
RTE 2022
Study on the technical conditions necessary for a power system with a High Share of Renewables in France Towards 2050, under the RCP 4.5 trajectory from IPCC 5th report.
@@
Shift 2020
The Shift Project published a plan for transforming French economy in 2022 for each sector of the economy. A preliminary report was published in end 2020 and this is the one we review here.
@@
France 2072 2018
This research work models two contrasting lifestyles for France in 2072,to answer the question: To what extent do lifestyles influence the energy system’s capacity to achieve carbon neutrality?; And amongs other, is a digital world compatible with the need to decarbonize the energy system, as usually thought?
@@
D&A 2022
This Digitalization & Anthropocene research study presents the past, present, and future of digitalization. It presents three illustrative future pathways that span the possibility space for digitalization and decarbonization in the Anthropocene. The authors conclude by identifying leverage points that shift human–digital– Earth interactions toward sustainability
@@
CNIL 2021
This study proposes new narratives and imaginaries regarding the protection of personal data in 2030 to help thinking regulatory practices. Throughout the process, highlights have been made on the use of technology in different social groups and at different times of digital life, as well as the resulting risks to individual and collective freedoms.
@@
Digit. Challenge 2022
The Université de Montréal and Espace pour la vie propose Chemins de transition (”Paths for transition”). This project combines a participative approach with major systemic challenges on a provincial scale. It combines different methodologies (backcasting, change-oriented approaches, transition management, etc.) and addresses 3 challenges: food, digital, territory. We here focus on digital challenge only.
@@ Appendix B Scenario variables
Figure 3: Mindmap of variables describing the scenarios

### v3

@@ Appendix A Summary of studies
Ademe 2022 This study from French Agency for Ecological Transition ADEME proposes five coherent and contrasting paths leading France towards carbon neutrality by 2050.
@@
negaWatt 2021 Scenarios for France made by the négaWatt association to achieve carbon neutrality by 2050 while reducing the extraction of raw materials. This scenario is also compatible with the -55% greenhouse gas target set at European level for 2030.
@@
EU green deal 2019 This communication sets out the strategy to transform the EU into a fair and prosperous society, with a modern, resource-efficient and competitive economy where there are no net emissions of greenhouse gases in 2050 and where economic growth is decoupled from resource use.
@@
Eionet 2022 This study produces a set of imaginaries offering engaging, plausible and contrasting images of what a sustainable Europe could look like in 2050. The imaginaries were developed through a participatory process, involving European Environment Agency staff, experts from the Eionet group on foresight, and external stakeholders.
@@
Arup 2019 Study made by an international multi-disciplinary team of consultants, designers and scientists. In the study, Climate considerations come third, subordinate to economic development and societal wellbeing. Scenarios are narratives and accompanied by a story told from the perspective of a person within that world, a timeline of events, and the achievement of the 17 UN SDGs.
@@
DDC 2020 The Danish Design Center (DDC) designed the Living Futures: Scenario Kit for understanding, discussing, and shaping the future. The kit is built around a set of four alternative futures set in the year 2050 explored through narrated stories by fictive people living in them.
@@
. SNBC 2020 The National Low Carbon Strategy (SNBC) describes a roadmap for France on how to steer its climate change mitigation policy. The SNBC is based on a reference scenario developed through a modelling exercise also used in the Multi-annual Energy Programming. The scenario details the public policy measures, in addition to those already in place, which will allow France to meet its short-, medium- and long-term climate and energy objectives as best it can.
@@
RTE 2022 Study on the technical conditions necessary for a power system with a High Share of Renewables in France Towards 2050, under the RCP 4.5 trajectory from IPCC 5th report.
@@
Shift 2020 The Shift Project published a plan for transforming French economy in 2022 for each sector of the economy. A preliminary report was published in end 2020 and this is the one we review here.
@@
France 2072 2018 This research work models two contrasting lifestyles for France in 2072,to answer the question: To what extent do lifestyles influence the energy system’s capacity to achieve carbon neutrality?; And amongs other, is a digital world compatible with the need to decarbonize the energy system, as usually thought?
@@
D&A 2022 This Digitalization & Anthropocene research study presents the past, present, and future of digitalization. It presents three illustrative future pathways that span the possibility space for digitalization and decarbonization in the Anthropocene. The authors conclude by identifying leverage points that shift human–digital– Earth interactions toward sustainability
@@
CNIL 2021 This study proposes new narratives and imaginaries regarding the protection of personal data in 2030 to help thinking regulatory practices. Throughout the process, highlights have been made on the use of technology in different social groups and at different times of digital life, as well as the resulting risks to individual and collective freedoms.
@@
Digit. Challenge 2022 The Université de Montréal and Espace pour la vie propose Chemins de transition (”Paths for transition”). This project combines a participative approach with major systemic challenges on a provincial scale. It combines different methodologies (backcasting, change-oriented approaches, transition management, etc.) and addresses 3 challenges: food, digital, territory. We here focus on digital challenge only.
@@ Appendix B Scenario variables
Variables Goal and design Project goal Process design Scenario content Societal variables Demography Lifestyles Economy Governance Non-digital tech. ICT variables Infrastructures Usage & Society Levers from equipment Applications Technologies •  Inclusion of norms
•  Vantage point
•  Subject
•  Time scale
•  Spatial scales
•  Nature of the data
•  Method of data collection
•  Nature of the resources
•  Nature of institutional conditions
•  Temporal nature
•  Nature of the variables
•  Nature of the dynamics
•  Level of deviation
•  Level of integration
• Importance of the state
• Fragmentation of the world
• Inequalities
• Economic growth
• Decoupling
• Financing transition
• Re-industrialization
• Sobriety/frugality
• Urban development
• Work
• Education
• Type of farming
• Resource circularity and availability
• Decarbonation
• Teleworking • Transportation • Industry • Surveillance • (Precise) medecine • Agriculture • Biology • Administration • Economy / Currency • Art and Leisure • Energy • Buildings • Urban planing • Monitoring • Optimisation of other sectors • Sensibilisation on environment • Communication • Education • e-commerce • Hacking and malicious use • Data analysis • AI • Cloud/edge comp. • Automation/robotics • IoT • Drones • Satellite images • Virtual/Mixed Reality • 3D printing • Open data • Environmental impact assessment (LCA) • Mobile networks • Low-tech • Evolution of usages
• Sobriety
• Digital Social Equity
• Digital behavior changes
• Data management and policies
• Data flow
• Number of user terminals
• Number of Network equipment
• Number of datacenters
• Datacenters location
• Cloud
• Data center consumption • Ecodesign of equipment
• Equipment lifetime
• Energy efficiency
Figure 3. Mindmap of variables describing the scenarios